Find all values of the parameter for which each of the roots of the equation x^3 - 3x - a = 0 not less than (a/10) + (6/5). In response, write down the sum of the integer values of the parameter a such that their modulus is less than 61.
1
Expert's answer
2020-12-21T16:41:45-0500

Solution-1:

1) We should simply the given equation:


"x(x^2-3)=a"


in this case, we assume that:

"x_1=a\;"

"x^2-3=1" from this we obtain the remaining roots of the equation:

"x_2=2\;;\;x_3=-2"


Let`s back to the question:

I)

"\frac{a}{10}+\frac{6}{5}\geq-2"


"a\geq-32"


II)

"a+|-2|+2\lt61"

"a\lt57"


Answer:"a\;\epsilon [-32,57)"



Solution-2:

1) We should simply the given equation:


"x(x^2-3)=a"


in this case, we assume that:

"x_1=1\;"

"x^2-3=a" from this we obtain the remaining roots of the equation:

"x_2=\sqrt{\smash[b]{a+3}}\;;\;x_3=-\sqrt{\smash[b]{a+3}}"


Let`s back to the question:

I)

"\frac{a}{10}+\frac{6}{5}\geq-\sqrt{\smash[b]{a+3}}"


"a\;\epsilon[-\infty,\infty]" but the domain of "x_2\;,x_3" should be "a\geq\;-3"


II)

"1+|-\sqrt{\smash[b]{a+3}}|+|\sqrt{\smash[b]{a+3}}|\lt61"

"a\lt897"


Answer:

"a\;\epsilon [-3,897)"